OpenStudy (patrickweewoos):

Paul’s Plumbing charges a flat fee of $90 plus $40/h for a plumbing job. Perfect Pipes charges a flat fee of $60 plus $55/h for a plumbing job. Write expressions to represent the total costs of a plumbing job for each company. Let h represent the time, in hours, of a plumbing job. Is the cost of a 3-hour-long plumbing job the same at both companies? Write and use an equation to find out.

OpenStudy (patrickweewoos):

@nincompoop

OpenStudy (patrickweewoos):

@ganeshie8

OpenStudy (patrickweewoos):

@Jhannybean

OpenStudy (patrickweewoos):

@Data_LG2

OpenStudy (anonymous):

a) paul's plumbing: c= 90 + 40h perfect pipes: c= 60 + 50h

OpenStudy (anonymous):

b) 90+40h h= 3 40 x 3 = 120 + 90= $210 perfect pipes: 60+50h h=3 50x3= 150+60=$210

OpenStudy (patrickweewoos):

I have to follow a rubric can you help? I dont understand it

OpenStudy (patrickweewoos):

ALGEBRA I CONSTRUCTED-RESPONSE SCORING GUIDELINES 16 – 20 points (possibly 17 – 20) Student response exhibits a comprehensive understanding of the mathematical concepts and procedures required of the assignment. Student response provides the correct answer(s) with clear and complete processes used and explained correctly. There are no major omissions in work or explanation distracting from exhibiting understanding. 11 - 15 points (possibly 14 - 16) Student response demonstrates a general understanding of the mathematical concepts and procedures required of the assignment. Student response provides mostly correct answer(s) and complete explanation exhibiting a general understanding of the concepts required. Only minor errors or omissions are evident. 6 – 10 points (possibly 11 - 13) Student response exhibits a partial understanding of the mathematical concepts and procedures required of the assignment. Student response provides partially correct answer and some explanation exhibiting little understanding of the concepts required. Major errors or omissions are evident. 0 – 5 points (possibly 0 – 10) Student response exhibits a no or minimal understanding of the mathematical concepts and procedures required of the assignment. Student response provides no or incorrect answer with no or incorrect explanation exhibiting minimal understanding of the concepts required.

OpenStudy (anonymous):

wats dat

OpenStudy (patrickweewoos):

A rubric

OpenStudy (anonymous):

oh of wat

OpenStudy (patrickweewoos):

I dont understand it.....I have to follow it and its of the problem i posted!

OpenStudy (anonymous):

ohh ok give me a sec

OpenStudy (patrickweewoos):

ok

OpenStudy (anonymous):

wait so the answer i gave u wasnt right or nah ???

OpenStudy (patrickweewoos):

Well..We have to follow the rubric for the problem. I just dont want to get it wrong

OpenStudy (anonymous):

ohh ok i understand

OpenStudy (patrickweewoos):

ok

OpenStudy (anonymous):

u mind if i tag someone and they help a lil more im a lil stomped

OpenStudy (anonymous):

@YK.PAPI answers are right

OpenStudy (anonymous):

thank you @peachpi

OpenStudy (welshfella):

YK.PAPI's answers are all correct but some added explanations might give you some extra points. First explain what your variables represent:- 'Let c = the cost of the job, and h = the number of hours spent doing the job.' and for Part (b) begin with 'Let h = 3 hours ' and at the end 'The above results show that the cost of a 3 hour job is the same for both companies'.

imqwerty (imqwerty):

ok 1st lets find out how much fee will paul take :) ok